# Adaptive Bounding Box Uncertainties via Two-Step Conformal Prediction

### v1

@@ 0.A.2 Conformal class thresholding (ClassThr)
The procedure to produce class label predictions sets satisfying a coverage guarantee via conformal class thresholding (ClassThr) [50] is outlined in Algorithm 2. Note that we directly sketch the algorithm for its class-conditional variation which satisfies Equation 4.
@@
14:For a new test sample $(X_{n+1},Y_{n+1})$, a valid conformal prediction set for $X_{n+1}$ is given by $\hat{C}_{L}(X_{n+1})=\{y\in\mathcal{Y}:\hat{\pi}_{y}(X_{n+1})\geq 1-\hat{q}_{L}^{y}\},$ where $\hat{q}_{L}^{y}\in Q$. Validity refers to satisfying sample coverage with probability $(1-\alpha_{L})$ per class, i.e., the guarantee in Equation 4. We refer to Sadinle et al. [50] for a proof.
@@
Since Algorithm 2 is used to obtain prediction sets for class labels, we now define a calibration sample $(X_{i},Y_{i})$ as consisting of an input image $X_{i}$ and the class label of an existing object in that image, i.e., we have $Y_{i}:=l_{i}$ following the object parametrization of subsection 2.2. Similarly, we have $\hat{C}_{L}(X_{n+1})\subseteq\mathcal{Y}:=\{1,\dots,K\}$ for a given classification task with $K$ classes.
@@ 0.A.3 Multiple testing problem
We show that running Algorithm 1 in parallel $m$ times with a global miscoverage rate $\alpha_{B}$ leads to the need for a corrected miscoverage rate on the individual level. This can be interpreted as a multiple testing correction, utilizing the hypothesis testing perspective motivated in subsection 4.1.
@@
Starting from Equation 3 for a fixed class $y$, we then have that
@@
Figure 5: Exact marginal coverage distributions for our three considered datasets (COCO, Cityscapes, BDD100k) on the basis of their calibration set sizes. For each distribution we additionally mark the $1\%$ and $99\%$ quantiles, as well as the desired target coverage. Left: For the conformal class label prediction sets on the basis of a target miscoverage rate $\alpha_{L}=0.01$. Right: For the conformal box prediction intervals on the basis of a target miscoverage rate $\alpha_{B}=0.1$. Obtained empirical coverage levels across experiments fall within reasonable regions of the analytically derived coverage distributions.
@@ 0.A.4 Multiple testing correction via $\max$-rank
The suggested correction is an adaptation of the Westfall & Young [67] min-P permutation correction. Intuitively, $\max$-rank collapses multiple testing dimensions into a single hypothesis test, using a composite empirical null distribution consisting of the maximum rank statistics across testing dimensions. The computed quantile over these maximum rank statistics will then ensure global coverage by design, since it controls coverage over ‘worst-case’ ranks.
Timans et al. [58] crucially relate their procedure to the hypothesis testing perspective of conformal prediction. In such a setting, each testing dimension’s empirical null distribution corresponds to the ranks of values in the computed sets of nonconformity scores $S$, i.e., for each of our box coordinates. It is shown that $\max$-rank satisfies two key requirements for its integration into a CP procedure – it preserves both exchangeability and validity. These conditions are demonstrated for the two key operations of the procedure, which are 1) operating in the rank domain of nonconformity scores, and 2) applying the $\max$-operator over these ranks. We refer to their work for further details.
@@ 0.A.5 Bounding box quantile selection
In subsection 5.2 we argue that a straight-forward quantile selection strategy is $\hat{q}_{B}^{k}=\max\{\hat{q}_{B}^{k,y}\}_{y\in\hat{C}_{L}(X_{n+1})}$, where $\hat{q}_{B}^{k,y}$ is the conformal quantile of the $k$-th box coordinate for class $y$. Then, for the $k$-th coordinate of a test sample whose true class label is indeed $y$ – which we ensure is contained in $\hat{C}_{L}(X_{n+1})$ with probability $(1-\alpha_{L})\approx 1$ – we have that $\hat{q}_{B}^{k}\geq\hat{q}_{B}^{k,y}$ and thus $\mathbb{P}(s(\hat{c}^{k}_{n+1},c^{k}_{n+1})\leq\hat{q}_{B}^{k})\geq\mathbb{P}(s(\hat{c}^{k}_{n+1},c^{k}_{n+1})\leq\hat{q}_{B}^{k,y})$. That is, using $\hat{q}_{B}^{k}$ as a thresholding quantile ensures a higher chance of score conformity, and thus is more likely to decide on inclusion into the prediction interval, at the cost of potential overcoverage.
From a hypothesis testing perspective, rather than using the p-value $p(c^{k,y})$, we base our testing decision on inclusion using $p(c^{k})=\max\{p(c^{k,y})\}_{y\in\hat{C}_{L}(X_{n+1})}$, itself a valid but conservative p-value [63]. Similarly, we then have that $p(c^{k})\geq p(c^{k,y})$ and thus $\mathbb{P}(p(c^{k})>\alpha_{B})\geq\mathbb{P}(p(c^{k,y})>\alpha_{B})$, producing a higher chance of inclusion. The coverage difference between testing with $p(c^{k})$ and $p(c^{k,y})$ then marks the incurred cost in terms of inefficiency, i.e., overly large prediction interval width.
@@ 0.A.6 Empirical coverage distribution
As mentioned in subsection 6.1, empirical coverage $Cov$ is a random quantity parametrized by a coverage distribution. For a specific randomly sampled calibration set $\mathcal{D}_{cal}$, achieved coverage may deviate from the target coverage level $(1-\alpha)$ based on factors such as calibration set size. Specifically, the exact analytical form of the coverage distribution is given by
@@
In Figure 5 we display the exact coverage distributions for our three considered datasets on the basis of their calibration set sizes for both label (left) and box (right) target coverage rates. We additionally mark the $1\%$ and $99\%$ quantiles of respective distributions. We see that obtainable empirical coverage levels are in line with our experiments, where results are aggregated over multiple trials of calibration set sampling. For example, in Figure 4, where we shade the coverage distribution between the above quantiles in grey ([Uncaptioned image]), obtained coverage levels using the Oracle method are within reasonable coverage bounds, even if on the lower end. Similarly, the magnitude of variations in coverage displayed in the violin plots of Figure 3 relates directly to respective calibration set sizes. This also holds for any additional results in Appendix 0.C.
@@ 0.B.3 Dataset splits and class mappings
We display the distribution of objects per class for the selected set of classes in Table 3. Objects are assigned to either calibration or test data based on the assignment to either split for the respective image they belong to. We randomly split the images according to the following calibration set sizes as a fraction of total available data for each dataset: 50% for COCO, 50% for Cityscapes and 70% for BDD100k.
@@
Table 4: Objects stratified by size as a fraction (in %) of total object counts (Table 3) for the selected set of common classes.
In Table 4 we additionally display object counts in % stratified by size, as measured via box surface area: small (area $\leq 32^{2}$), medium (area $\in(32^{2},96^{2}]$) and large (area $>96^{2}$). While Cityscapes contains virtually no small objects, we generally observe a mix of different object sizes for various classes.
@@
  • $\bullet$
    for Cityscapes, we map classes ‘pedestrian’ and ‘rider’ to class ‘person’;
  • $\bullet$
    for BDD100k, we map classes ‘person’ and ‘rider’ to class ‘person’.
@@ 0.B.4 Model details and parameter settings
    The score parameter removes any box proposals that receive a model confidence score below a specified threshold. We fix this value at a confidence of 0.5.
@@ 0.B.5 Prior work comparison
We compare our methods to the conformal bounding box approaches presented in Andéol et al. [1] – an extension of [14] – which we consider closest prior work. As mentioned in section 3, their work proposes conformal scoring methods designed for one-sided, outer prediction interval construction. We thus modify our conformal box coordinate methods (Box-Std and Box-Ens) to produce similar one-sided intervals for comparison purposes. We do not consider modifying Box-CQR, because it is not straightforward how a one-sided version of its conformal scores should be constructed.
In addition, Andéol et al. [1] only consider a single object class, thus bypassing the intricacy of capturing uncertainty in the class label predictions, as our two-step approach does. Since our comparison relates to differences in the box interval constructions only, we evaluate both their and our modified approaches across the set of classes using the correct class quantiles only (i.e., via the Oracle).
Let us once again consider the 2D bounding box setting with explicit coordinate tuples $Y=(c^{1},c^{2},c^{3},c^{4})$. To maintain notational consistency with Andéol et al. [1], we rename the coordinates by grouping them into two pairs denoting box corners as $Y:=(x_{0},y_{0},x_{1},y_{1})$. We then compare to their following proposed conformal scoring methods:
@@
    and obtain outer prediction interval coordinates as
@@
    where $\hat{q}_{B}$ are the respective coordinate-level conformal quantiles at coverage level $(1-\alpha^{\prime}_{B})$, and $\alpha^{\prime}_{B}=\alpha_{B}/4$ is the Bonferroni correction for target coverage $(1-\alpha_{B})$.
@@
    where $\hat{w}=\hat{x}_{1}-\hat{x}_{0}$ and $\hat{h}=\hat{y}_{1}-\hat{y}_{0}$ are the predicted box width and height respectively. Outer intervals are then constructed as
@@
    where once again $\alpha^{\prime}_{B}=\alpha_{B}/4$ is corrected via Bonferroni.
@@
    AddMax, MultMax. Instead of using a Bonferroni correction, Andéol et al. [1] also suggest taking a $\max$-operation over coordinate scores in Equation 12 and Equation 14 respectively, resulting in a set of ‘maximal’ scores. A conformal quantile $\hat{q}_{B}$ is then computed directly at target coverage over these scores and used for each coordinate, alleviating the need for further correction. The idea is close to the $\max$-rank approach [58], but instead operates directly in the domain of scores. This can be problematic due to an improper influence of score magnitudes across dimensions. Resulting outer interval coordinates are then constructed as in Equation 13 (AddMax) and Equation 15 (MultMax) simply by replacing the quantiles.
@@
    Box-Std. We use the conformal scoring approach from Equation 12 in conjunction with the $\max$-rank multiple testing correction. This is the equivalent of a one-sided signed version of the regression rsiduals emplyed by Box-Std.
@@
    Box-Ens. We create a one-sided version of Box-Ens by using the conformal scores from Equation 14, but using ensemble-fused coordinate predictions and normalizing by the obtained ensemble uncertainties $\hat{\sigma}$.
@@
    Box-Mult. We additionally consider using the conformal scores from Equation 14 directly with a $\max$-rank correction, a combination approach that can be related to the normalized scores of Box-Ens.
Box stretch metric. We implement the proposed ‘box stretch’ evaluation metric from Andéol et al. [1] that assesses the additional box surface area incurred by CP. Formally, we denote the metric as
@@
where $\mathcal{A}(\cdot)$ is the computed surface area of the bounding box formed by the respective input, i.e., the conformal outer interval bounds, and the predicted bounding box coordinates. Ideally, we desire $Stretch$ to be close to 1.0.
Mean prediction interval width. The $MPIW$ metric is formally defined for two-sided intervals only. In order to also allow comparison using $MPIW$ alongside $Stretch$, we transform all of the above one-sided interval constructions into two-sided intervals. We do so by considering the distance of the outer interval coordinates to the box center, i.e., we place a lower interval bound at the predicted box center coordinates. Note that we also do the same for our own initially two-sided methods to allow for a fair comparison.
@@ 0.C.1 Effect of calibration on Naive and ClassThr
We aim to better understand the strong empirical coverage results observed in Figure 4 for the Naive label prediction set baseline. Since Naive is theoretically motivated from a model calibration perspective, we empirically assess its sensitivity to changes in calibration, and compare results to ClassThr. We refer to Guo et al. [19] for more details on calibration and related concepts.
Specifically, we proxy model calibration via top-class calibration, which can be evaluated using the Expected Calibration Error (ECE) metric (see Figure 6). We then influence model calibration via temperature scaling [19] – a simple logit-scaling approach – to induce regimes of model miscalibration via over- and underconfidence. We find that our object detector (temperature $T=1.0$) is generally slightly overconfident, and optimal confidence in terms of ECE lies at $T^{*}=1.3$ (Figure 6, top). We then perform our two-step conformal approach using Box-Std, and compare ClassThr to the use of Naive for multiple severities of over- and underconfidence (Figure 6, bottom). Both methods are evaluated in regards to label coverage $(1-\alpha_{L})$ (left axis) and mean set size (right axis).
We find that as miscalibration through overconfidence increases (bottom left), Naive gradually fails to satisfy target coverage of 99 %. In contrast, ClassThr is able to continuously provide desired coverage levels, at the cost of increasing set sizes for high miscalibration. As miscalibration through underconfidence increases (bottom right), the set sizes of Naive explode, tending towards the full domain (mean set size $\sim 80$). In contrast, ClassThr provides tighter coverage levels and maintains reasonable set sizes. We conclude that while Naive provides good empirical results under relatively accurate model calibration (as seen in Figure 4) it is quite sensitive to miscalibration. In contrast, ClassThr seems to provide a more robust safety assurance even under settings of over- and underconfidence, at the risk of potentially large set sizes.

### v2

@@ 0.A.2 Conformal class thresholding (ClassThr)
The procedure to produce class label predictions sets satisfying a coverage guarantee via conformal class thresholding (ClassThr) [52] is outlined in Algorithm 2. Note that we directly sketch the algorithm for its class-conditional variation which satisfies Eq. 4. We also invert the scores used by Sadinle et al. [52] to better reflect the notion of nonconformity, while being equivalent in procedure.
@@
14:For a new test sample $(X_{n+1},Y_{n+1})$, a valid prediction set for $X_{n+1}$ is given by $\hat{C}_{L}(X_{n+1})=\{y\in\mathcal{Y}:\hat{\pi}_{y}(X_{n+1})\geq 1-\hat{q}_{L}^{y}\},$ where $\hat{q}_{L}^{y}\in Q$. Validity refers to satisfying sample coverage with probability $(1-\alpha_{L})$ per class, i.e., the guarantee in Eq. 4. We refer to Sadinle et al. [52] for a proof.
@@
Since Algorithm 2 is used to obtain prediction sets for class labels, we now define a calibration sample $(X_{i},Y_{i})$ as consisting of an input image $X_{i}$ and the class label of an existing object in that image, i.e., we have $Y_{i}:=l_{i}$ following the object parametrization of § 2.2. Similarly, we have $\hat{C}_{L}(X_{n+1})\subseteq\mathcal{Y}:=\{1,\dots,K\}$ for a given classification task with $K$ classes.
@@ 0.A.3 Multiple testing problem
Conformal prediction can also be interpreted from a hypothesis testing perspective [65, 55]. A null hypothesis for the $k$-th test sample coordinate is formed as $H_{0}:c^{k}_{n+1}=c^{k}$ for some candidate value $c^{k}\in\mathbb{R}$. Leveraging the nonconformity scores of the calibration samples as an empirical null distribution, a p-value $p(c^{k})$ is computed as the fraction of samples which conform worse than $s(\hat{c}^{k}_{n+1},c^{k})$  [^5] A valid p-value $P$ is defined as $\mathbb{P}(P\leq\alpha)\leq\alpha\quad\forall\alpha\in(0,1).$. The value $c^{k}$ is included in the prediction set — i.e., we cannot reject $H_{0}$ — if $p(c^{k})>\alpha_{B}$. Conformalizing a set of $m$ coordinates separately and thus running Algorithm 1 in parallel $m$ times with a global miscoverage rate $\alpha_{B}$ leads to the need for a corrected miscoverage rate on the individual level, as we show next.
@@
Starting from Eq. 3 for a fixed class $y$, we then have that
@@ 0.A.4 Multiple testing correction via $\max$-rank
The suggested correction is an adaptation of the Westfall & Young [70] min-P permutation correction. Intuitively, $\max$-rank collapses multiple testing dimensions into a single hypothesis test, using a composite empirical null distribution consisting of the maximum rank statistics across testing dimensions. The computed quantile over these maximum rank statistics will then ensure global coverage by design, since it controls coverage over ‘worst-case’ ranks. Timans et al. [61] crucially relate their procedure to the hypothesis testing perspective of conformal prediction. In such a setting, each testing dimension’s empirical null distribution corresponds to the ranks of values in the computed sets of nonconformity scores $S$, i.e., for each of our box coordinates. It is shown that $\max$-rank satisfies two key requirements for its integration into a CP procedure – it preserves both exchangeability and validity. These conditions are demonstrated for the two key operations of the procedure, which are 1) operating in the rank domain of nonconformity scores, and 2) applying the $\max$-operator over these ranks. We refer to their work for further details.
@@ 0.A.5 Bounding box quantile selection
In § 5.2 we argue that a straight-forward quantile selection strategy is $\hat{q}_{B}^{k}=\max\{\hat{q}_{B}^{k,y}\}_{y\in\hat{C}_{L}(X_{n+1})}$, where $\hat{q}_{B}^{k,y}$ is the conformal quantile of the $k$-th box coordinate for class $y$. Then, for the $k$-th coordinate of a test sample whose true class label is indeed $y$ – which we ensure is contained in $\hat{C}_{L}(X_{n+1})$ with probability $(1-\alpha_{L})\approx 1$ – we have that $\hat{q}_{B}^{k}\geq\hat{q}_{B}^{k,y}$ and thus $\mathbb{P}(s(\hat{c}^{k}_{n+1},c^{k}_{n+1})\leq\hat{q}_{B}^{k})\geq\mathbb{P}(s(\hat{c}^{k}_{n+1},c^{k}_{n+1})\leq\hat{q}_{B}^{k,y})$. That is, using $\hat{q}_{B}^{k}$ as a thresholding quantile ensures a higher chance of score conformity, and thus is more likely to decide on inclusion into the prediction interval, at the cost of potential overcoverage.
From a hypothesis testing perspective, rather than using the p-value $p(c^{k,y})$, we base our testing decision on inclusion using $p(c^{k})=\max\{p(c^{k,y})\}_{y\in\hat{C}_{L}(X_{n+1})}$, itself a valid but conservative p-value [66]. Similarly, we then have that $p(c^{k})\geq p(c^{k,y})$ and thus $\mathbb{P}(p(c^{k})>\alpha_{B})\geq\mathbb{P}(p(c^{k,y})>\alpha_{B})$, producing a higher chance of inclusion. The coverage difference between testing with $p(c^{k})$ and $p(c^{k,y})$ then marks the incurred cost in terms of inefficiency, i.e., prediction interval width.
Figure 5: Exact marginal coverage distributions for our three considered datasets (COCO, Cityscapes, BDD100k) on the basis of their calibration set sizes. For each distribution we additionally mark the $1\%$ and $99\%$ quantiles, as well as the desired target coverage. Left: For the conformal class label prediction sets on the basis of a target miscoverage rate $\alpha_{L}=0.01$. Right: For the conformal box prediction intervals on the basis of a target miscoverage rate $\alpha_{B}=0.1$. Obtained empirical coverage levels across experiments fall within reasonable regions of the derived coverage distributions.
@@ 0.A.6 Empirical coverage distribution
As mentioned in § 6.1, empirical coverage $Cov$ is a random quantity parametrized by a coverage distribution. For a specific randomly sampled calibration set $\mathcal{D}_{cal}$, achieved coverage may deviate from the target coverage level $(1-\alpha)$ based on factors such as calibration set size. Specifically, the exact analytical form of the coverage distribution is given by
@@
In Fig. 5 we display the exact coverage distributions for our three considered datasets on the basis of their calibration set sizes for both label (left) and box (right) target coverage rates. We additionally mark the $1\%$ and $99\%$ quantiles of respective distributions. We see that obtainable empirical coverage levels are in line with our experiments, where results are aggregated over multiple trials of calibration set sampling. For example, in Fig. 4, where we shade the coverage distribution between the above quantiles in grey ([Uncaptioned image]), obtained coverage levels using the Oracle method are within reasonable coverage bounds, even if on the lower end. Similarly, the magnitude of variations in coverage displayed in the violin plots of Fig. 3 relates directly to respective calibration set sizes. This also holds for any additional results in Appendix 0.C.
@@ 0.B.3 Dataset splits and class mappings
We display the distribution of objects per class for the selected set of classes in Tab. 3. Objects are assigned to either calibration or test data based on the assignment to either split for the respective image they belong to. We randomly split the images according to the following calibration fractions of total available data for each dataset: 50% for COCO, 50% for Cityscapes and 70% for BDD100k.
@@
Table 4: Objects stratified by size as a fraction (in %) of total object counts (Tab. 3) for the selected set of common classes.
In Tab. 4 we additionally display object counts in % stratified by size, as measured via box surface area: small (area $\leq 32^{2}$), medium (area $\in(32^{2},96^{2}]$) and large (area $>96^{2}$). While Cityscapes contains virtually no small objects, we generally observe a mix of different object sizes for various classes.
@@ 0.B.4 Model details and parameter settings
    The score parameter removes box proposals that receive a model confidence score below a specified threshold. We fix this value at a confidence of 0.5.
@@ 0.B.5 Prior work comparison
We compare our methods to the conformal bounding box methods presented in Andéol et al. [1] – an extension of [15] – which we consider closest prior work. As mentioned in § 3, they propose conformal scoring functions designed for one-sided, outer prediction intervals. We thus modify our bounding box methods Box-Std and Box-Ens to produce similar one-sided intervals for comparison. We do not modify Box-CQR because it is unclear how a one-sided version of its scoring function should be constructed. Andéol et al. [1] only consider a single object class, thus bypassing the need to account for uncertainty in the class label predictions, as we do in § 5. Since our comparison relates to differences in the bounding box step only, we evaluate both their and our modified approaches across the set of classes using correct class quantiles (i.e., via the Oracle).
Let us reconsider the 2D bounding box setting with explicit coordinate tuples $(c^{1},c^{2},c^{3},c^{4})$. To maintain notational consistency with Andéol et al. [1], we rename the coordinates into two pairs denoting box corners as $Y:=(x_{0},y_{0},x_{1},y_{1})$. We then compare to the following proposed scoring functions:
@@
    and obtain outer prediction intervals as
@@
    where $\hat{q}_{B}$ are the respective coordinate-level conformal quantiles at coverage level $(1-\alpha^{\prime}_{B})$, and $\alpha^{\prime}_{B}=\alpha_{B}/4$ is the corrected Bonferroni coverage level.
@@
    where $\hat{w}=\hat{x}_{1}-\hat{x}_{0}$ and $\hat{h}=\hat{y}_{1}-\hat{y}_{0}$ are predictions for box width and box height. Outer box intervals are then constructed as
@@
    where once more $\alpha^{\prime}_{B}=\alpha_{B}/4$ is the Bonferroni correction.
@@
    AddMax, MultMax. Instead of a Bonferroni correction, Andéol et al. [1] also suggest a $\max$-operation over coordinate scores in Eq. 12 and Eq. 14 respectively, resulting in a set of ‘maximal’ scores. A conformal quantile $\hat{q}^{\max}_{B}$ is then computed directly at target coverage level over these scores and used across coordinates, alleviating the need for further correction. The principle is related to the $\max$-rank approach [61], but instead operates directly in the domain of scores, which can be problematic due to an improper influence of score magnitudes. Outer box intervals are then constructed as in Eq. 13 (AddMax) and Eq. 15 (MultMax) by replacing the quantiles with $\hat{q}^{\max}_{B}$.
@@
    Box-Std. We use the conformal scoring approach from Eq. 12 in conjunction with the $\max$-rank multiple testing correction. This is the equivalent of a one-sided signed version of the regression residuals employed by Box-Std.
@@
    Box-Ens. We create a one-sided version of Box-Ens by using the conformal scores from Eq. 14, but using ensemble-fused coordinate predictions and normalizing by the obtained ‘heuristic’ ensemble uncertainty $\hat{\sigma}$.
@@
    Box-Mult. We additionally consider using the conformal scores from Eq. 14 directly with a $\max$-rank correction, a combination approach that can be related to the normalized scores of Box-Ens.
Box stretch metric. We implement the proposed box stretch evaluation metric from Andéol et al. [1] that assesses the additional box surface area incurred by applying conformal prediction. Formally, we denote the metric as
@@
where $\mathcal{A}(\cdot)$ is the computed surface area of the bounding box formed by the respective input, i.e., the outer box interval bounds, and the predicted bounding box. Ideally, we desire $Stretch$ to be close to 1.0.
Mean prediction interval width. The $MPIW$ metric is formally defined for two-sided intervals only. In order to permit comparison using $MPIW$ alongside $Stretch$, we transform any of the above one-sided interval constructions into two-sided intervals. We do so by considering the distance of the outer interval coordinates to the box center, i.e., placing a lower interval bound at the predicted box center coordinates. Note that we also do the same for our own initially two-sided methods to allow for a fair comparison.
@@ 0.C.1 Effect of calibration on Naive and ClassThr
We aim to better understand the strong empirical coverage results observed in Fig. 4 for the Naive label prediction set baseline. Since Naive is theoretically motivated from a model calibration perspective, we empirically assess its sensitivity to changes in calibration, and compare results to ClassThr. We refer to Guo et al. [20] for more details on calibration and related concepts.
Specifically, we proxy model calibration via top-class calibration, which can be evaluated using the Expected Calibration Error (ECE) metric (see Fig. 6). We then influence model calibration via temperature scaling [20] – a simple logit-scaling approach – to induce regimes of model miscalibration via over- and underconfidence. We find that our object detector (temperature $T=1.0$) is generally slightly overconfident, and optimal confidence in terms of ECE lies at $T^{*}=1.3$ (Fig. 6, top). We then perform our two-step conformal approach using Box-Std, and compare ClassThr to the use of Naive for multiple severities of over- and underconfidence (Fig. 6, bottom). Both methods are evaluated in regards to label coverage $(1-\alpha_{L})$ (left axis) and mean set size (right axis).
We find that as miscalibration through overconfidence increases (bottom left), Naive gradually fails to satisfy target coverage of 99 %. In contrast, ClassThr is able to continuously provide desired coverage levels, at the cost of increasing set sizes for high miscalibration. As miscalibration through underconfidence increases (bottom right), the set sizes of Naive explode, tending towards the full domain (mean set size $\sim 80$). In contrast, ClassThr provides tighter coverage levels and maintains reasonable set sizes. We conclude that while Naive provides good empirical results under relatively accurate model calibration (as seen in Fig. 4) it is quite sensitive to miscalibration. In contrast, ClassThr seems to provide a more robust safety assurance even under settings of over- and underconfidence, at the cost of potentially larger set sizes.